Saturday, 22 August

Saturday, 22 August

Invited Lectures

\underline{14:00 – 14:45}

03 Vladimir Berkovich, Weizmann Institute of Science, Rehovot, Israel p-adic analytic spaces (I: p.48, electronic version) Chair: Norbert Schappacher

MA 005

07 James Greig Arthur, University of Toronto, Toronto, Canada Towards a stable trace formula (I: p.87) Chair: Michael Rapoport

MA 004

08 Kari Astala, University of Jyväskylä, Jyväskylä, Finland Analytic aspects of quasiconformality (I: p.97) Chair: Frederick W. Gehring

MA 001

09 Yuri Kifer, Hebrew University of Jerusalem, Jerusalem, Israel Random Dynamics and its Applications (I: p.114) Chair: Ludwig Arnold

H 1012

13 Béla Bollobás, University of Memphis, Memphis, USA and University of Cambridge, Cambridge, UK Hereditary properties of graphs: asymptotic enumeration, global structure, and colouring (I: p.153) Chair: Hans J. Prömel

H 2032

16 Thomas Yizhao Hou, California Institute of Technology, Pasadena, USA Numerical Study of Free Interface Problems Using Boundary Integral Methods (I: p.189) Chair: Vladimir V. Shaidurov

H 1028

\underline{15:00 – 15:45}

03 Pierre Colmez, CNRS, ENS, and Institut de Mathématiques de Jussieu, Paris, France (φ, Γ)-modules and Iwasawa theory of p-adic representations (I: p.49) Chair: Christophe Soulé

MA 005

07 Ivan V. Cherednik, The University of North Carolina at Chapel Hill, Chapel Hill, USA From double Hecke algebra to analysis (I: p.89) Chair: Olivier Mathieu

MA 004

08 Detlef H. Müller, Universität Kiel, Kiel, Germany L^{p}-functional calculus on Lie groups and estimates for wave equations (I: p.102) Chair: N.N.

MA 001

09 Curtis Tracy McMullen, Harvard University, Cambridge, USA Rigidity and inflexibility in conformal dynamics (I: p.119) Chair: Joerg Schmeling

H 1012

13 Harald Niederreiter, Austrian Academy of Sciences, Vienna, Austria Nets, (t, s)-sequences, and algebraic curves over finite fields with many rational points (I: p.159) Chair: Dieter Jungnickel

H 2032

16 Leslie Frederick Greengard, Courant Institute, New York University, New York, USA A new version of the fast Gauss transform (I: p.189) Chair: Gregory Beylkin

H 1028

\underline{16:15 – 17:00}

02 Eric Mark Friedlander, Northwestern University, Evanston, USA Geometry of Infinitesimal Group Schemes (I: p.36) Chair: Uwe Jannsen MA 005 04 & 11 Paul S. Aspinwall, Duke University, Durham, USA Geometry and String Duality (I: p.59) Chair: Pelham M.H. Wilson

H 1028

05 Dominic Joyce, Lincoln College, Oxford, UK Compact Riemannian manifolds with exceptional holonomy (I: p.73) Chair: Boris Dubrovin

MA 001

06 William Gerard Dwyer, University of Notre Dame, Notre Dame, USA Lie groups and p-compact groups (I: p.79) Chair: Albrecht Dold

MA 004

10 Gunther A. Uhlmann, University of Washington, Seattle, USA Inverse Boundary Value Problems for Partial Differential Equations (I: p.127) Chair: Bert-Wolfgang Schulze

H 1012

17 William R. Pulleyblank, IBM T.J. Watson Research Center, Yorktown Heights, USA Advances in Solving Airline Crew Pairing Optimization Problems (I: p.194) Chair: Rolf H. Möhring

H 2032

\underline{17:15 – 18:00}

!!!

Lecture by a Fields Medalist

– to be announced – Chair: N.N.

H 2032

02 Gunter Malle, Universität Heidelberg, Heidelberg, Germany Spetses (I: p.41) Chair: John S. Wilson

MA 005

05 Sylvestre Gallot, U.M.P.A., École Normale Supérieure de Lyon, Lyon, France Curvature-decreasing maps are volume-decreasing (I: p.71) Chair: Boris Dubrovin

MA 001

06 Mark Mahowald, Northwestern University, Evanston, USA Toward a global understanding of π_{∗}(S^{n}) (I: p.81) Chair: Albrecht Dold

MA 004

10 Mikhail Vasilevich Safonov, University of Minnesota, Minneapolis, USA Estimates near the boundary for solutions of second order parabolic equations (I: p.126) Chair: Heinz O. Cordes

H 1012

11 Pierre Collet, CNRS, École Polytechnique, Palaiseau, France Extended Dynamical Systems (I: p.130) Chair: Roberto Schonmann H 1028